# Reinforcement Learning with Hidden Markov Models for Discovering Decision-Making Dynamics

### v1

@@ 4.3 Brain-behavior Association
We examine whether abnormalities in reward processing may reflect underlying changes in the brain circuits that regulate reward-related behavior and learning. Specifically, we investigate whether dysfunctions in MDD-implicated brain circuitries affect learning abilities by an association analysis of subjects’ individual engagement scores and task functional magnetic resonance imaging (fMRI) measures of brain activation. We focus on fMRI measures in an emotional conflict task assessing amygdala-anterior cingulate (ACC) circuitry (Etkin et al., 2006; Fonzo et al., 2019). The emotional conflict task comprises four types of trials: congruent (C) trials, incongruent (I) trials, incongruent followed by incongruent (iI) trials, and incongruent followed by congruent (cI) trials.
@@
To investigate the brain-behavior association for each brain region of interest (ROI), the voxel-level activation conflicts (I-C) and activation conflict adaptations (iI - cI) within each ROI for each subject in the MDD group before treatment were computed. To investigate interregional interaction between ROIs, a psychophysiological interaction (PPI) analysis, as outlined in Etkin et al. (2006), provides voxel-wise conflict adaptation coupling between two ROIs. As a result, we considered six types of fMRI measures: (i) mean of activation conflicts within ROIs (I-C mean), (ii) mean of activation conflict adaptations within ROIs (iI-cI mean), (iii) mean of conflict adaptation coupling between two ROIs (PPI mean), (iv) standard deviation of activation conflicts within ROIs (I-C std), (v) standard deviation of activation conflict adaptations within ROIs (iI-cI std), and (vi) standard deviation of conflict adaptation coupling between two ROIs (PPI std). We utilize linear regressions to model the associations between individual engagement scores and each fMRI measure for all ROIs/interactions. Subsequently, we calculate the p-values for the regression coefficients and implement q-values for controlling the false discovery rate (FDR), as proposed by (Benjamini and Hochberg, 1995).
@@
The results are depicted in Figure 4. In Figure 4(a), each $q$-value undergoes a transformation using a negative $\log_{10}$ function, with the dashed line representing the FDR at $10\%$. Three ROIs and two ROI interactions were detected. A visualization illustrating the detected ROIs or interactions in the brain is presented in Figure 4(b), where the associations between engagement scores and the brain response to emotional conflict adaptation coupling of subgenual ACC ROI and left amygdala ROI appears to be the most significant. Other significant associations include the conflict coupling of subgenual ACC and right amygdala, the conflict within dorsal ACC and right insula, and the conflict adaptations within the subgenual ACC. In this analysis, all significant brain measures identified were the standard deviations of ROIs, rather than their mean effects. In addition, our results show a negative correlation between engagement scores and the standard deviations of all five identified ROIs or interactions. This suggests that an increased engagement in reward learning tasks corresponds to a decreased variability in brain activity during an emotional conflict task. Williams (2016) further indicated that the ACC—amygdala connectivity belongs to the negative affect circuit. Hence, MDD patients are more prone to lose concentration during the reward learning behavioral task, which could be due to an increase of fluctuation (or instability) of brain activities within the negative affect circuit.
@@ 5 Discussion
In this paper, we propose a RL-HMM method to characterize reward-based decision-making, incorporating the switching of learning strategies between two states: the ‘engaged’ state, where subjects make decisions based on the RL model, and the ‘lapse’ state, where they make random choices. We extend the RL model to accommodate a non-discrete state space and posit time-varying transition probabilities within the HMM. For parameter estimation, we employ the EM algorithm. In the E-step, we apply forward-backward inductive computation, and in the M-step, we implement a computationally efficient algorithm for one-step updates of all the parameters. Extensive simulation studies validate the finite-sample performance of our method.
@@
Applying our approach to the EMBARC, we discovered different phenomena from the findings in Huys et al. (2013) and Guo et al. (2023), where the MDD group had a lower reward sensitivity compared to the healthy controls. In contrast, our method did not identify a significant group difference in reward sensitivity. Instead, we discovered that the MDD group spent a larger proportion of time using the ‘lapse’ strategy compared to the healthy group. This suggests an alternative mechanism of the reward processing abnormalities in MDD. We identified a significant correlation between the individual engagement score and a clinician-assessed measure of concentration difficulty. This finding suggests that tendencies towards disengagement may be linked with a loss of concentration ability in MDD. Moreover, our analysis revealed that the ‘lapse’ strategy requires less time for individuals to respond than the ‘engaged’ strategy in PRT. Using additional brain fMRI data within the same study, we found that the individual engagement scores computed from RL-HMM are associated with the variability of brain activities in the negative affect circuitry under an emotional conflict task (Etkin et al., 2006). This brain-behavior association study suggests that a higher heterogeneity in brain activities during emotional conflict is associated with lower engagement in reward learning tasks.
@@
Our methods are sufficiently general to analyze other types of behavioral tasks. Some extensions include to introduce covariates to model between-subjects heterogeneity in learning processes and jointly model PRT with other phenotypes including brain measures and clinical outcomes.

### v2

@@ 4.3 Brain-behavior Association
We examine whether abnormalities in reward processing may reflect underlying changes in the brain circuits that regulate reward-related behavior and learning. Specifically, we investigate whether dysfunctions in MDD-implicated brain circuitries affect learning abilities by an association analysis of subjects’ individual engagement scores and task functional magnetic resonance imaging (fMRI) measures of brain activation. We focus on fMRI measures in an emotional conflict task assessing amygdala-anterior cingulate (ACC) circuitry (Etkin et al.,, 2006; Fonzo et al.,, 2019). The emotional conflict task comprises four types of trials: congruent (C) trials, incongruent (I) trials, incongruent followed by incongruent (iI) trials, and incongruent followed by congruent (cI) trials.
@@
To investigate the brain-behavior association for each brain region of interest (ROI), the voxel-level activation conflicts (I-C) and activation conflict adaptations (iI - cI) within each ROI for each subject in the MDD group before treatment were computed. To investigate interregional interaction between ROIs, a psychophysiological interaction (PPI) analysis, as outlined in Etkin et al., (2006), provides voxel-wise conflict adaptation coupling between two ROIs. As a result, we considered six types of fMRI measures: (i) mean of activation conflicts within ROIs (I-C mean), (ii) mean of activation conflict adaptations within ROIs (iI-cI mean), (iii) mean of conflict adaptation coupling between two ROIs (PPI mean), (iv) standard deviation of activation conflicts within ROIs (I-C std), (v) standard deviation of activation conflict adaptations within ROIs (iI-cI std), and (vi) standard deviation of conflict adaptation coupling between two ROIs (PPI std). We utilize linear regressions to model the associations between individual engagement scores and each fMRI measure for all ROIs/interactions. Subsequently, we calculate the p-values for the regression coefficients and implement q-values for controlling the false discovery rate (FDR), as proposed by (Benjamini and Hochberg,, 1995).
@@
The results are depicted in Figure 3. In Figure 3(a), each $q$-value undergoes a transformation using a negative $\log_{10}$ function, with the dashed line representing the FDR at $10\%$. Three ROIs and two ROI interactions were detected. A visualization illustrating the detected ROIs or interactions in the brain is presented in Figure 3(b), where the associations between engagement scores and the brain response to emotional conflict adaptation coupling of subgenual ACC ROI and left amygdala ROI appears to be the most significant. Other significant associations include the conflict coupling of subgenual ACC and right amygdala, the conflict within dorsal ACC and right insula, and the conflict adaptations within the subgenual ACC. In this analysis, all significant brain measures identified were the standard deviations of ROIs, rather than their mean effects. In addition, our results show a negative correlation between engagement scores and the standard deviations of all five identified ROIs or interactions. This suggests that an increased engagement in reward learning tasks corresponds to a decreased variability in brain activity during an emotional conflict task. Williams, (2016) further indicated that the ACC—amygdala connectivity belongs to the negative affect circuit. Hence, MDD patients are more prone to lose concentration during the reward learning behavioral task, which could be due to an increase of fluctuation (or instability) of brain activities within the negative affect circuit.
@@ 5 Discussion
In this paper, we propose a RL-HMM method to characterize reward-based decision-making, incorporating the switching of decision-making strategies between two states: the ‘engaged’ state, where subjects make decisions based on the RL model, and the ‘lapse’ state, where they make random choices. We extend the RL model to accommodate a non-discrete state space and posit time-varying transition probabilities within the HMM. For parameter estimation, we employ the EM algorithm. In the E-step, we apply forward-backward inductive computation, and in the M-step, we implement a computationally efficient algorithm for one-step updates of all the parameters. Extensive simulation studies validate the finite-sample performance of our method.
@@
Applying our approach to the EMBARC, we discovered different phenomena from the findings in Huys et al., (2013) and Guo et al., (2024), where the MDD group had a lower reward sensitivity compared to the healthy controls. In contrast, our method did not identify a significant group difference in reward sensitivity. Instead, we discovered that the MDD group spent a larger proportion of time using the ‘lapse’ strategy compared to the healthy group. This suggests an alternative mechanism of the reward processing abnormalities in MDD. We identified a significant correlation between the individual engagement score and a clinician-assessed measure of concentration difficulty. This finding suggests that tendencies towards disengagement may be linked with a loss of concentration ability in MDD. Moreover, our analysis revealed that the ‘lapse’ strategy requires less time for individuals to respond than the ‘engaged’ strategy in PRT. Using additional brain fMRI data within the same study, we found that the individual engagement scores computed from RL-HMM are associated with the variability of brain activities in the negative affect circuitry under an emotional conflict task (Etkin et al.,, 2006). This brain-behavior association study suggests that a higher heterogeneity in brain activities during emotional conflict is associated with lower engagement in reward learning tasks.
@@
In PRT, The individual engagement score serves as a crucial behavioral marker. However, our current work does not accommodate heterogeneity beyond the aspects of the pattern of individual engagement probabilities during tasks. Some extensions include introducing covariates to model between-subjects heterogeneity in RL parameters (such as learning rate, reward sensitivity, and initial coefficients of the expected reward) and transition probabilities in the HMM. Our methods are sufficiently general to analyze other types of behavioral tasks. Another extension involves jointly modeling PRT with other phenotypes, including brain measures and clinical outcomes.
Recall that we assume subjects make decisions independently of how they learn from rewards. In psychology, humans may learn from rewards without conscious awareness; this phenomenon is termed implicit learning (Frensch and Rünger,, 2003). It suggests that individuals can unconsciously detect patterns and learn from the outcomes of their actions over time. In the reinforcement learning (RL) literature, another phenomenon, exploration versus exploitation, may help in understanding our assumption: Exploration in RL involves testing new actions to discover potential rewards, while exploitation entails selecting actions known to yield high rewards. Even within an exploration strategy, reward learning still occurs (Sutton and Barto,, 2018). A generalization of our proposed model is that we assume reward learning occurs in different decision-making strategies, and the learning behavior—encoded by RL parameters such as learning rate and reward sensitivity—might vary depending on the strategy. For example, individuals in the ‘lapse’ state may learn more slowly or be less sensitive to reward.
6 Supporting Materials
Supplementary materials containing additional algorithm and simulation details are available online at http://biostatistics.oxfordjournals.org. The R code for the simulation studies is available on GitHub at https://github.com/xingcheg/RL-HMM.
Acknowledgments
This research is supported by U.S. NIH grants MH123487, NS073671 and GM124104.